OpenStudy (anonymous):

Find the slope and y intercept of the graph y=5-x/10

OpenStudy (***[isuru]***):

when a graph is present by the format y = mx + c m is the gradient or the slope c is the intercept y = 5 - x/10 = -x/10 + 5